If you used 61.5mL of 3.80 M KOH to titrate 15.0 ML of H2SO4, what was the concentration of the H2SO4?.
1
Expert's answer
2020-03-19T02:35:37-0400

Solution:

Consider the option of complete neutralization.

The balanced chemical equation of the reaction of complete neutralization:

2KOH + H2SO4 = K2SO4 + 2H2O.

At the equivalence point:

n(KOH) / 2 = n(H2SO4);

C(KOH) * V(KOH) = 2 * C(H2SO4) * V(H2SO4);

The molarity of the H2SO4 is calculated as follows:

C(H2SO4) = [C(KOH) * V(KOH)] / [2 * V(H2SO4)] = [3.80 M * 0.0615 L] / [2 * 0.0150 L] = 7.79 M

C(H2SO4) = 7.79 M


Answer: 7.79 M was the concentration of the H2SO4.